# Understanding Prompt Tuning and In-Context Learning via Meta-Learning

### v1

@@ 4.2 Results
Fig. 1 shows that both, a Transformer and an LSTM, pretrained on Random Coins, can be Soft Prompted to be Bayes-optimal on a Single Coin (with bias $0.2$). Other prefix-tuning methods, including exhaustive search over all hard token sequences of length $6$ fail to reach Bayes-optimality. Despite all soft prefixes being off-distribution inputs, internal dynamics remain stable (see Fig. 2), and prediction generalize well for ‘SoftPF’ and most of the weight-tuning methods far beyond the tuning sequence length $N_{\text{tune}}=50$ (see Fig. 1). The results illustrate how prefixes can be used to steer a (meta-learned) Bayesian predictor via manipulating its internal state. Further, they also show that off-distribution inputs can be particularly effective; more effective than even the best possible hard token sequence of the same length.
@@
Figure 4: Untrained Transformer tuned to the Two-Coin mixture (left) and to Random Coins (right); tuning sequence length ${N_{\text{tune}}=50}$. In both cases, Soft Prompting is the only effective prefix-tuning method. It nearly reaches Bayes-optimality (‘TargetBayes’, which is a Laplace predictor on Random Coins). Performance degrades rapidly after the tuning sequence length. Full regret curves (and LSTM results) in Fig. A8 and Fig. A9. Among the weight-tuning methods, LoRA is very effective.
@@
Fig. 3 empirically demonstrates the theoretical shortcoming of prefix-tuning discussed in Section 3: prompt tuning to a mixture of two coins is not possible if the pretraining distribution is uniform random coins. While Soft Prompting, being the strongest prefix-tuning method, leads to performance gains compared to the untuned pretrained predictor, it is not enough to reach ‘TargetBayes’ on the Two-Coin Mixture. As expected, some weight-tuning methods can lead to that level of performance, at the cost of permanently altering the pretrained predictor. While soft prefixes are strictly speaking not covered by the Bayesian theory (because they are off-distribution inputs exploiting the circuitry of the particular pretrained network), these results highlight the importance of the Bayesian view in practice: in the absence of theoretical understanding it might have been quite puzzling why one can optimally prefix prompt for a single coin but not a mixture of two coins. We confirm that the result is not an artifact of limited soft prefix length—see Fig. A10 for control experiment with $L=25$.
@@
Fig. 4 shows that it is possible to Soft Prompt an untrained Transformer quite well to the Two-Coin Mixture and Random Coins as target tasks, meaning that relatively complex in-context algorithms are easily available in the untrained net. There is still a gap to ‘TargetBayes’ performance though. Results in the Appendix (Figs. A8 and A9) show that Soft Prompting the untrained LSTM has very little effect, indicating a fundamental difference between the Transformer and the LSTM in this regard (which behave very similarly on our experiments when pretrained). See also Zhong and Andreas (2024), who tune untrained Transformers to algorithmic tasks via embedding- or unembedding-tuning, or both. They find that, tuning both the embedding and unembedding is important on their tasks. If our results qualitatively hold at their tasks, then Soft Prompting of untrained networks should be only slightly worse, and LoRA should perform even better than Un+Embedding tuning. Increasing the soft prefix length to $L=25$ does not significantly improve performance, see Fig. A10.
@@ 5 Discussion
Extrapolating our results to frontier-model scale, soft prefix tuning (specifically Soft Prompting) should be superior to tuning sequences of hard tokens—if a sufficiently large fine-tuning set is available, soft prompt tuning should beat prompt engineering, and other methods of hard token optimization such as PromptBreeder (Fernando et al., 2023). Similarly, instead of conditioning on a large set of in-context examples for imitation learning (e.g., Ruoss et al. (2025); Paglieri et al. (2025)), it may be beneficial to distill these examples into a more effective tuned soft prefix. The superiority of Soft Prompting over other soft prefix-tuning methods may be due to the much higher dimensionality of “embeddings” compared to the input space in our experiments. This is typically reversed in frontier models. Finally, our experiments raise the question: why prompt (-tune) at all, when weight-based methods, such as LoRA, are equally or more effective and do not suffer from the theoretical limitations pointed out? First, weight-tuning permanently alters a network and would lead to performance decreases on the pretraining distribution (which can be overcome by storing the set of original weights). Additionally, comparisons between in-context and in-weight learning at LLM scale find that in-weight learning can sometimes be very limited and generalize poorly (Lampinen et al., 2025; Chan et al., 2022b). Since prompting fundamentally builds on a network’s in-context adaptation mechanisms, it may be the case that prompt-tuning works better than weight-tuning in cases where in-context learning generalizes better than in-weight learning. Another interesting future research question is whether tuned (soft) prefixes transfer between different models (perhaps with additional regularization)—if true, the prefix-tuning cost would only have to be spent once, compared to weight-tuning methods that need to be run for every network to fine-tune. Please see our discussion of additional related work in Appendix D, where we discuss a number of previous works that investigate in-context learning under a Bayesian and/or meta-learning lens, such as Xie et al. (2021), Kirsch et al. (2022), Lampinen et al. (2024), and Elmoznino et al. (2024). To conclude, the Bayesian view on prompt tuning, which arises from analyzing memory-based meta-learning, provides a conceptual understanding that leads to a formal characterization of prompting and some of its limitations. We have shown that this holds in practice, for both hard token prefix-tuning, but also when optimizing soft prefixes—a setting not fully covered by the theory.
Acknowledgments.
@@ Appendix D Additional Related Work
In-context learning has been studied extensively in the recent literature. In many cases, it specifically refers to a particular type of supervised few-shot learning. In contrast, Lampinen et al. (2024) (among others) argue that a whole number of LLM in-context abilities can be unified as in-context learning in a wider sense: “[…] we suggest that any distribution of sequences in which context non-trivially decreases loss on subsequent predictions can be interpreted as eliciting a kind of in-context learning. We suggest that this perspective helps to unify the broad set of in-context abilities that language models exhibit.”. This is in line with our view on in-context learning, and we argue that the Bayesian perspective that arises from analyzing memory-based meta-learning provides the unifying theoretical framework. Lampinen et al. (2024) also discuss memory-based meta-learning as the underlying factor. We present this connection in more formal detail (see Section 2), and use it to drive the design of our experiments. Bayesian inference has also been put forward as an explanation for the mechanism that drives in-context learning in Xie et al. (2021); Müller et al. (2022); Genewein et al. (2023); Binz et al. (2024); Wang et al. (2023); Panwar et al. (2024).
@@
To adapt (large) pretrained models, many fine-tuning methods, with a number of variations each, have been proposed in the recent literature. See Han et al. (2024) for a review of parameter-efficient fine-tuning methods, such as, soft prompting (Lester et al., 2021), prefix prompting (Li and Liang, 2021), or LoRA, Hu et al. (2021). Whether a method falls under prefix-tuning or weight-tuning may not always be immediately obvious, since tunable inputs also appear as indirectly tunable parameters in query and key matrices of transformers. Careful analysis reveals though, that tunable parameters resulting from prefix-tuning methods are more constrained compared to weight based tuning. Petrov et al. (2024) perform such an analysis and find: “[…] while techniques like prompting, in-context learning, soft prompting, and prefix-tuning can effectively elicit skills present in the pretrained model, they may not be able to learn novel tasks that require new attention patterns.”. This empirical finding is in line with one of the theoretical negative results for prompting that we point out in Section 3. The difference between in-context learning (which is the mechanism that prompt tuning exploits) and in-weight learning has been studied in a series of works (Lampinen et al., 2025; Agarwal et al., 2024; Chan et al., 2022b) that find that in-context learning can typically generalize more flexibly than weight-tuning, and that the underlying mechanisms are quite different (e.g., rule-based vs. exemplar-based generalization), and may result from different neural circuits that compete during training (Singh et al., 2023, 2025), as well as different underlying properties of the data distribution (Chan et al., 2025). Kirsch et al. (2022) study general-purpose in-context learning via meta-learning and investigate the factors that lead models to generalize (via meta-learned in-context algorithms) as opposed to memorization. An interesting approach presented in Bornschein et al. (2023) is to switch from in-context to in-weight learning as soon as the available examples allow it (in terms of number of samples and statistical properties w.r.t. the pretrained predictor), which is determined via prequential evaluation.
@@ Appendix E Theoretical limits of prompting Bayesian predictors
[…]
@@
One can ask whether a larger (countable) class of distributions =${\mathcal{M}=\{\nu\}}$ always allows for optimal prefix-tuning: Let the mixture ${\xi(x):=\sum_{\nu\in\mathcal{M}}\nu(x)\tilde{w}(\nu)}$ with some prior ${\tilde{w}(\nu)>0}$. Indeed, for suitable $\mathcal{M}$ it holds that: for every computable (Bernoulli) prior $w$ there exists some $y\in\mathbb{B}^{*}$ such that $\xi(x|y)=\zeta_{w}(x)$ for all $x\in\mathbb{B}^{*}$. That is, one can prompt $\xi$ such that its predictive distribution becomes any desired Bernoulli mixture. The construction is quite artificial though: we require that the model class is such that the prefix is interpreted as a program that explicitly computes the desired target prior. Formally, let ${\nu_{p}(x_{1:n}):=[\![x_{<\ell}=p]\!]\zeta_{w}(x_{\ell:n})}$, where ${p\in\mathbb{B}^{*}}$ is a prefix program computing prior $w()$ and $\ell-1=$length$(p)$ (i.e. ${\mathcal{P}:=\{p:p~{}\text{computes some}~{}w\}}$ is a prefix-free set). Let =${\mathcal{M}=\{\nu_{p}:p\in\mathcal{P}\}}$. Then it is easy to see that ${\xi(x_{\ell:n}|x_{<\ell})=\zeta_{w}(x_{\ell:n})}$, where $x_{<\ell}\in\mathcal{P}$ computes prior $w$.
This class is very artificial, but it shows that prefix-tuning for general Bernoulli mixtures is possible in principle. No special property of $\zeta_{w}$ was used, so the above tuning construction works for arbitrary countable base class =$\mathcal{B}=\{\zeta_{i}\}$.
@@
To overcome the need that the prefix is a program that computes the prior, another construction is possible, where the prompt is a number of increasingly longer sequences of samples from the target distribution. Let =${\mathcal{B}=\{\zeta\}}$ be a countable class of target distributions (e.g. Bernoulli mixtures). Let ${1=i_{0}<i_{1}<i_{2}<...}$ be a sequence of temporal boundaries with increasing segment lengths ${\delta_{\kappa}:=i_{\kappa}-i_{\kappa-1}\to\infty}$ for ${\kappa\to\infty}$, e.g. $\delta_{\kappa}=\kappa$ or $\delta_{\kappa}=2^{\kappa}$. Define the product distribution as ${\nu_{\zeta}(x_{1:n}):=\prod_{\kappa=1}^{m}\zeta(x_{i_{\kappa-1}:i_{\kappa}-1})}$ for $n=i_{m}-1$, and for other $n$ by marginalization. Let ${\xi(x):=\sum_{\zeta\in\mathcal{B}}\nu_{\zeta}(x)\tilde{w}(\zeta)}$ be a Bayes mixture with some prior $\tilde{w}>0$. Then by (Hutter, 2005, Sec.3.7.1) or Blackwell and Dubins (1962) for any $h<\infty$ we have ${\xi(x_{\ell:\ell+h-1}|x_{<\ell})\to\nu_{\zeta}(x_{\ell:\ell+h-1}|x_{<\ell})}$ a.s. if ${x_{1:\infty}\sim\nu_{\zeta}}$. If ${i_{m}-h\geq\ell=i_{m-1}}$ for some $m$, then ${\nu_{\zeta}(x_{\ell:\ell+h-1}|x_{<\ell})=\zeta(x_{\ell:\ell+h-1})}$. The condition is satisfied for infinitely many $\ell$. Hence for any $\zeta\in\mathcal{B}$ and any $h\in\mathbb{N}$ there exists a prompt $x_{<\ell}$ such that ${\xi(x_{\ell:\ell+h-1}|x_{<\ell})\approx\zeta(x_{\ell:\ell+h-1})}$ for all $x_{\ell:\ell+h-1}\in\mathbb{B}^{h}$, and the approximation error can be made arbitrarily small by suitably large $\ell$.
@@
In expectation, extra information decreases entropy ($H(X|Y)\leq H(X)$), but specific information can increase or decrease entropy ($H(X|Y=y)\gtrless H(X)$). In our sequential context this means that $H(\cdot|X_{<\ell})\leq H(\cdot)$, where $\cdot$ can be $X_{\ell:\infty}$ or $X_{\ell:\ell+h-1}$ or $\theta$. This means under some ergodicity assumptions for large $\ell$, if $x_{<\ell}\sim P=\xi$, then likely $H(\cdot|x_{<\ell})\lesssim H(\cdot)$, i.e. typical prompts narrow the task distribution. In the Beta-Bernoulli case we even have $H(\theta|x_{<\ell})\to 0$ if $x_{1:\infty}\sim\xi$ whatever the prior $w(\theta)$, which implies that the posterior converges to a $\delta$-peak. Conversely if we want the posterior $w(\theta|x_{<\ell})$ to be broader (have higher entropy) than the prior $w(\theta)$, we need an atypical prompt $x_{<\ell}$. For instance, for a Beta-Bernoulli with prior $w(\theta)=\varepsilon\delta(\theta)+(1-\varepsilon)\delta(\theta-\frac{1}{2})$ and small $\varepsilon$, the entropy $H(\theta|x_{<\ell})$ increases with $\ell$ for small $\ell$ for the atypical prompt $x_{<\ell}=0^{\ell-1}$.

### v2

@@ 4.2 Results
Fig.˜1 shows that both, a Transformer and an LSTM, pretrained on Random Coins, can be Soft Prompted to be Bayes-optimal on a Single Coin (with bias $0.2$). Other prefix-tuning methods, including exhaustive search over all hard token sequences of length $6$ fail to reach Bayes-optimality. Despite all soft prefixes being off-distribution inputs, internal dynamics remain stable (see Fig.˜2), and prediction generalize well for ‘SoftPF’ and most of the weight-tuning methods far beyond the tuning sequence length $N_{\text{tune}}=50$ (see Fig.˜1). The results illustrate how prefixes can be used to steer a (meta-learned) Bayesian predictor via manipulating its internal state. Further, they also show that off-distribution inputs can be particularly effective; more effective than even the best possible hard token sequence of the same length.
Note that for all prefix-tuning methods the dimensionality of $\mathcal{S}$ is the same as the one-hot token alphabet (i.e., $2$-dimensional for coin-flip tasks), except for ‘SoftPT’ where the prefix embeddings of dimensionality $128$ are tuned. These additional degrees of freedom are the main source of superior performance in our experiments, and we demonstrate that the advantage largely disappears when reducing the embedding dimensionality to $4$ in Appendix˜I. Since LLMs typically have a larger input- than embedding dimensionality, tuning inputs (‘RealPT’) may be as efficient, or even more efficient, compared to tuning embeddings (‘SoftPT’) for LLMs.
@@
Figure 4: Untrained Transformer tuned to the Two-Coin mixture (left) and to Random Coins (right); tuning sequence length $N_{\text{tune}}=50$}. In both cases, Soft Prompting is the only effective prefix-tuning method. It nearly reaches Bayes-optimality (‘TargetBayes’, which is a Laplace predictor on Random Coins). Performance degrades rapidly after the tuning sequence length. Full regret curves (and LSTM results) in Fig.˜A8 and Fig.˜A9. Among the weight-tuning methods, LoRA is very effective.
@@
Fig.˜3 empirically demonstrates the theoretical shortcoming of prefix-tuning discussed in Section˜3: prompt tuning to a mixture of two coins is not possible if the pretraining distribution is uniform random coins. While Soft Prompting, being the strongest prefix-tuning method, leads to performance gains compared to the untuned pretrained predictor, it is not enough to reach ‘TargetBayes’ on the Two-Coin Mixture. As expected, some weight-tuning methods can lead to that level of performance, at the cost of permanently altering the pretrained predictor. While soft prefixes are strictly speaking not covered by the Bayesian theory (because they are off-distribution inputs exploiting the circuitry of the particular pretrained network), these results highlight the importance of the Bayesian view in practice: in the absence of theoretical understanding it might have been quite puzzling why one can optimally prefix prompt for a single coin but not a mixture of two coins. We confirm that the result is not an artifact of limited soft prefix length—see Fig.˜A10 for a control experiment with $L=25$, and Fig.˜A12 for results with larger networks.
@@
Fig.˜4 shows that it is possible to Soft Prompt an untrained Transformer quite well to the Two-Coin Mixture and Random Coins as target tasks, meaning that relatively complex in-context algorithms are easily available in the untrained net. There is still a gap to ‘TargetBayes’ performance though. Results in the Appendix (Figs.˜A8 and A9) show that Soft Prompting the untrained LSTM has very little effect, indicating a fundamental difference between the Transformer and the LSTM in this regard (which behave very similarly on our experiments when pretrained). See also Zhong and Andreas (2024), who tune untrained Transformers to algorithmic tasks via embedding- or unembedding-tuning, or both. They find that, tuning both the embedding and unembedding is important on their tasks. If our results qualitatively hold at their tasks, then Soft Prompting of untrained networks should be only slightly worse, and LoRA should perform even better than Un+Embedding tuning. As previously, note that the superiority of ‘SoftPT’ in our setting is largely explained by having a much higher embedding dimensionality compared to the input dimensionality (which is typically reverse at LLM scale). Increasing the soft prefix length to $L=25$ does not significantly improve performance, see Fig.˜A10.
@@ 5 Discussion
While we have laid important fundamental groundwork in our current study, extrapolating our findings to modern frontier model (and data) scale is not straightforward. The theoretical findings we presented, including the limitations of prefix tuning, hold at any scale, but it is likely that additional practical issues arise at large scale that are not captured by our current experiments. It is thus hard to predict the relevance and impact of our fundamental results on today’s frontier-model practice. For instance, one of our main results is that optimal prompting to a single target task is possible, whereas it is not for a mixture of tasks. We are confident that this holds even at frontier model scale (based on the theory), but it is unclear and highly non-trivial what constitutes a task for a LLM, and accordingly, whether this is a severe limitation or not. In our experiments, a task is simply an unobserved variable in a two-level hierarchical statistical model. At LLM scale, the structure is vastly more complex, with many more hierarchical levels, and potentially other statistical structures at play. The next step would be to carefully design data generators that are closer to natural language data, but still fully understood and well controllable, akin to the data generators used in (Allen-Zhu, 2024), and run our experiments of tuning to single tasks vs. tuning to mixtures of tasks at scale. With these caveats in mind, cautiously extrapolating our findings to frontier model scale, raises some questions for investigation, which we now discuss.
Given our results, soft prefix tuning should be superior to tuning sequences of hard tokens—if a sufficiently large fine-tuning set is available, soft prompt tuning should beat prompt engineering, and other methods of hard token optimization such as PromptBreeder (Fernando et al., 2023) [^3] Though some hard-token tuning methods may have the advantage of resulting in more interpretable prompts.. Similarly, instead of conditioning on a large set of in-context examples for imitation learning (e.g., Ruoss et al. (2025); Paglieri et al. (2025)), it may be beneficial to distill these examples into a more effective tuned soft prefix. The superiority of Soft Prompting over other soft prefix-tuning methods is largely due to the much higher dimensionality of “embeddings” compared to the input space in our experiments (see control experiments with embedding dimensionality $4$ in Appendix˜I). This is typically reversed in frontier models, which could mean that soft input tuning is more effective than embedding tuning.
@@
Finally, our experiments raise the question: why prompt (-tune) at all, when weight-based methods, such as LoRA, are equally or more effective and do not suffer from the theoretical limitations pointed out? First, weight-tuning permanently alters a network and would lead to performance decreases on the pretraining distribution (which can be overcome by storing the set of original weights). Additionally, comparisons between in-context and in-weight learning at LLM scale find that in-weight learning can sometimes be very limited and generalize poorly (Lampinen et al., 2025; Chan et al., 2022b). Since prompting fundamentally builds on a network’s in-context adaptation mechanisms, it may be the case that prompt-tuning works better than weight-tuning in cases where in-context learning generalizes better than in-weight learning. An interesting future research question is whether tuned (soft) prefixes transfer between different models (perhaps with additional regularization)—if true, the prefix-tuning cost would only have to be spent once, compared to weight-tuning methods that need to be run for every network to fine-tune. Please see our discussion of additional related work in Appendix˜D, where we discuss a number of previous works that investigate in-context learning under a Bayesian and/or meta-learning lens, such as Xie et al. (2021), Kirsch et al. (2022), Lampinen et al. (2024), and Elmoznino et al. (2024).
To conclude, the Bayesian view on prompt tuning, which arises from analyzing memory-based meta-learning, provides a conceptual understanding that leads to a formal characterization of prompting and some of its fundamental limitations. We have shown that these limitations hold in practice, for both hard token prefix-tuning, but also when optimizing soft prefixes—a setting not fully covered by the theory (which does not consider non-token inputs). The code to reproduce all our experiments and figures is available at: https://github.com/google-deepmind/thunnini.
Acknowledgments and Disclosure of Funding
@@ Appendix D Additional Related Work
In-context learning has been studied extensively in the recent literature. In many cases, it specifically refers to a particular type of supervised few-shot learning. In contrast, Lampinen et al. (2024) (among others) argue that a whole number of LLM in-context abilities can be unified as in-context learning in a wider sense: “[…] we suggest that any distribution of sequences in which context non-trivially decreases loss on subsequent predictions can be interpreted as eliciting a kind of in-context learning. We suggest that this perspective helps to unify the broad set of in-context abilities that language models exhibit.”. This is in line with our view on in-context learning, and we argue that the Bayesian perspective that arises from analyzing memory-based meta-learning provides the unifying theoretical framework. Lampinen et al. (2024) also discuss memory-based meta-learning as the underlying factor. We present this connection in more formal detail (see Section˜2), and use it to drive the design of our experiments. Bayesian inference has also been put forward as an explanation for the mechanism that drives in-context learning in Xie et al. (2021); Müller et al. (2022); Genewein et al. (2023); Binz et al. (2024); Wang et al. (2023); Panwar et al. (2024).
@@
To adapt (large) pretrained models, many fine-tuning methods, with a number of variations each, have been proposed in the recent literature. See Han et al. (2024) for a review of parameter-efficient fine-tuning methods, such as, soft prompting (Lester et al., 2021), prefix prompting (Li and Liang, 2021), or LoRA, Hu et al. (2021). Whether a method falls under prefix-tuning or weight-tuning may not always be immediately obvious, since tunable inputs also appear as indirectly tunable parameters in query and key matrices of transformers. Careful analysis reveals though, that tunable parameters resulting from prefix-tuning methods are more constrained compared to weight based tuning. Petrov et al. (2024) perform such an analysis and find: “[…] while techniques like prompting, in-context learning, soft prompting, and prefix-tuning can effectively elicit skills present in the pretrained model, they may not be able to learn novel tasks that require new attention patterns.”. This empirical finding is in line with one of the theoretical negative results for prompting that we point out in Section˜3. The difference between in-context learning (which is the mechanism that prompt tuning exploits) and in-weight learning has been studied in a series of works (Lampinen et al., 2025; Agarwal et al., 2024; Chan et al., 2022b) that find that in-context learning can typically generalize more flexibly than weight-tuning, and that the underlying mechanisms are quite different (e.g., rule-based vs. exemplar-based generalization), and may result from different neural circuits that compete during training (Singh et al., 2023, 2025), as well as different underlying properties of the data distribution (Chan et al., 2025). Kirsch et al. (2022) study general-purpose in-context learning via meta-learning and investigate the factors that lead models to generalize (via meta-learned in-context algorithms) as opposed to memorization. An interesting approach presented in Bornschein et al. (2023) is to switch from in-context to in-weight learning as soon as the available examples allow it (in terms of number of samples and statistical properties w.r.t. the pretrained predictor), which is determined via prequential evaluation.
@@ Appendix E Theoretical limits of prompting Bayesian predictors
Let $\mu_{\theta}(x_{1:n}):=\theta^{k}(1-\theta)^{n-k}$ with $k_{n}:=x_{1}+...+x_{n}$ and $x_{t}\in\mathbb{B}:=\{0,1\}$ be a Bernoulli($\theta$) process with prior (density) $w(\theta)$ for $\theta\in[0;1]$. Then the posterior is $w(\theta|x_{1:n})=\mu_{\theta}(x_{1:n})w(\theta)/\zeta_{w}(x_{1:n})$, where normalizer = mixture = evidence $\zeta_{w}(x_{1:n}):=\int_{0}^{1}\mu_{\theta}(x_{1:n})w(\theta)d\theta$.
If $w(\theta)$ is a Beta prior (e.g. uniform), then the posterior is also Beta-distributed with variance $\text{Var}[w(\cdot|x_{1:n})]\leq k(n-k)/n^{3}\leq 1/4n\to 0$ for $n\to\infty$. Hence the posterior of $w$ “converges” to a $\delta$-peak whatever $x_{1:\infty}$.
More generally, for any prior such that $w$ and sequence $x_{1:\infty}$ for which all limit points of $\hat{\theta}_{n}:=k_{n}/n$ are in the support of $w$, we have $\text{Var}[w(\cdot|x_{1:n})]\to 0$. Note this holds even if $k_{n}/n$ itself does not converge.
@@
One can ask whether a larger (countable) class of distributions $\mathcal{M}=\{\nu\}$} always allows for optimal prefix-tuning: Let the mixture $\xi(x):=\sum_{\nu\in\mathcal{M}}\nu(x)\tilde{w}(\nu)$} with some prior $\tilde{w}(\nu)>0$}. Indeed, for suitable $\mathcal{M}$ it holds that: for every computable (Bernoulli) prior $w$ there exists some $y\in\mathbb{B}^{*}$ such that $\xi(x|y)=\zeta_{w}(x)$ for all $x\in\mathbb{B}^{*}$. That is, one can prompt $\xi$ such that its predictive distribution becomes any desired Bernoulli mixture. The construction is quite artificial though: we require that the model class is such that the prefix is interpreted as a program that explicitly computes the desired target prior. Formally, let $\nu_{p}(x_{1:n}):=[\![x_{<\ell}=p]\!]\zeta_{w}(x_{\ell:n})$}, where $p\in\mathbb{B}^{*}$} is a prefix program computing prior $w()$ and $\ell-1=$length(p)(p) (i.e. $\mathcal{P}:=\{p:p~\text{computes some}~w\}$} is a prefix-free set). Let $\mathcal{M}=\{\nu_{p}:p\in\mathcal{P}\}$}. Then it is easy to see that $\xi(x_{\ell:n}|x_{<\ell})=\zeta_{w}(x_{\ell:n})$}, where $x_{<\ell}\in\mathcal{P}$ computes prior $w$.
This class is very artificial, but it shows that prefix-tuning for general Bernoulli mixtures is possible in principle. No special property of $\zeta_{w}$ was used, so the above tuning construction works for arbitrary countable base class $\mathcal{B}=\{\zeta_{i}\}$.
@@
To overcome the need that the prefix is a program that computes the prior, another construction is possible, where the prompt is a number of increasingly longer sequences of samples from the target distribution. Let $\mathcal{B}=\{\zeta\}$} be a countable class of target distributions (e.g. Bernoulli mixtures). Let $1=i_{0}<i_{1}<i_{2}<$...} be a sequence of temporal boundaries with increasing segment lengths $\delta_{\kappa}:=i_{\kappa}-i_{\kappa-1}\to\infty$} for $\kappa\to\infty$}, e.g. $\delta_{\kappa}=\kappa$ or $\delta_{\kappa}=2^{\kappa}$. Define the product distribution as $\nu_{\zeta}(x_{1:n}):=\prod_{\kappa=1}^{m}\zeta(x_{i_{\kappa-1}:i_{\kappa}-1})$} for $n=i_{m}-1$, and for other $n$ by marginalization. Let $\xi(x):=\sum_{\zeta\in\mathcal{B}}\nu_{\zeta}(x)\tilde{w}(\zeta)$} be a Bayes mixture with some prior $\tilde{w}>0$. Then by (Hutter, 2005, Sec.3.7.1) or Blackwell and Dubins (1962) for any $h<\infty$ we have $\xi(x_{\ell:\ell+h-1}|x_{<\ell})\to\nu_{\zeta}(x_{\ell:\ell+h-1}|x_{<\ell})$} a.s. if $x_{1:\infty}\sim\nu_{\zeta}$}. If $i_{m}-h\geq\ell=i_{m-1}$} for some $m$, then $\nu_{\zeta}(x_{\ell:\ell+h-1}|x_{<\ell})=\zeta(x_{\ell:\ell+h-1})$}. The condition is satisfied for infinitely many $\ell$. Hence for any $\zeta\in\mathcal{B}$ and any $h\in\mathbb{N}$ there exists a prompt $x_{<\ell}$ such that $\xi(x_{\ell:\ell+h-1}|x_{<\ell})\approx\zeta(x_{\ell:\ell+h-1})$} for all $x_{\ell:\ell+h-1}\in\mathbb{B}^{h}$, and the approximation error can be made arbitrarily small by suitably large $\ell$.
@@
In expectation, extra information decreases entropy ($H(X|Y)\leq H(X)$), but specific information can increase or decrease entropy ($H(X|Y=y)\gtrless H(X)$). In our sequential context this means that $H(\cdot|X_{<\ell})\leq H(\cdot)$, where $\cdot$ can be $X_{\ell:\infty}$ or $X_{\ell:\ell+h-1}$ or $\theta$. This means under some ergodicity assumptions for large $\ell$, if $x_{<\ell}\sim P=\xi$, then likely $H(\cdot|x_{<\ell})\lesssim H(\cdot)$, i.e. typical prompts narrow the task distribution. In the Beta-Bernoulli case we even have $H(\theta|x_{<\ell})\to$-$\infty$ if $x_{1:\infty}\sim\xi$ whatever the prior $w(\theta)$, and the posterior converges to a $\delta$-peak. Conversely if we want the posterior $w(\theta|x_{<\ell})$ to be broader (have higher entropy) than the prior $w(\theta)$, we need an atypical prompt $x_{<\ell}$. For instance, for a Beta-Bernoulli with prior $w(\theta)=\varepsilon\delta(\theta)+(1-\varepsilon)\delta(\theta-\frac{1}{2})$ and small $\varepsilon$, the entropy $H(\theta|x_{<\ell})$ increases with $\ell$ for small $\ell$ for the atypical prompt $x_{<\ell}=0^{\ell-1}$.